This problem is the difference of rational expressions.
y / 6x - x+4 / 3x
or y divided by 6x minus x+4 divided by 3x.
Please show me how to solve completely. Thanks in advance!

Answer by edjones(7569)

y/6x - (x+4)/3x
=(y-(2(x+4)))/6x LCD=6x
=(y-2x-8)/6x
.
Ed